# Online Dense Point Tracking with Streaming Memory

### v1

@@ Abstract
Dense point tracking is a challenging task requiring the continuous tracking of every point in the initial frame throughout a substantial portion of a video, even in the presence of occlusions. Traditional methods use optical flow models to directly estimate long-range motion, but they often suffer from appearance drifting without considering temporal consistency. Recent point tracking algorithms usually depend on sliding windows for indirect information propagation from the first frame to the current one, which is slow and less effective for long-range tracking. To account for temporal consistency and enable efficient information propagation, we present a lightweight and fast model with Streaming memory for dense POint Tracking and online video processing. The SPOT framework features three core components: a customized memory reading module for feature enhancement, a sensory memory for short-term motion dynamics modeling, and a visibility-guided splatting module for accurate information propagation. This combination enables SPOT to perform dense point tracking with state-of-the-art accuracy on the CVO benchmark, as well as comparable or superior performance to offline models on sparse tracking benchmarks such as TAP-Vid and RoboTAP. Notably, SPOT with 10$\times$ smaller parameter numbers operates at least 2$\times$ faster than previous state-of-the-art models while maintaining the best performance on CVO. We will release the models and codes at: https://github.com/DQiaole/SPOT.
@@ 3.2 Feature Enhancement with Memory Reading
Specifically, the original feature $\mathbf{F}_{t}$ is projected through a linear projection layer for generating the query feature $q$. Then a dot-product similarity calculation is performed against the memory key followed by a softmax for normalization. This normalized score is used to aggregate the memory value to obtain the readout features $\mathbf{F}_{t}^{\prime}$. The process can be formulated as:
$\mathbf{F}_{t}^{\prime}=\mathrm{Softmax}(1/\sqrt{D_{k}}\times q\times k^{T})\times v,$ (1)
@@
Yet, we would like to point out that it is inappropriate to use the readout features $\mathbf{F}_{t}^{\prime}$ for flow regression directly. As our memory value is obtained through splatting the reference feature $\mathbf{F}_{1}$ with the previously predicted flow (which will be introduced in Sec. 3.4), there are many blank hole artifacts [1, 31] due to disocclusion [27]. These disocclusion artifacts will hamper the network training and degrade the performance of flow regression. Inspired by the typically used inpainting technique for hole filling [1], we further propose a fusion layer for artifacts remedy. Fortunately, we find that one single convolution layer performs well enough, and we “inpaint” the feature through the original feature $\mathbf{F}_{t}$:
$\hat{\mathbf{F}}_{t}=\mathbf{F}_{t}+\mathrm{Conv}(\mathbf{F}_{t}\oplus\mathbf{F}_{t}^{\prime}),$ (2)
where $\oplus$ denotes concatenation and $\hat{\mathbf{F}}_{t}$ is the final enhanced feature that is ready for the next stage of flow regression.
@@ 3.3 Flow Decoding and Sensory Memory
Figure 3: Overview of flow and sensory memory update. Motion feature is obtained through flow lookup. We update the flow with sensory memory, and update the sensory memory with the final motion feature. For the sake of simplicity, visibility prediction and context feature are omitted here.
With the memory-enhanced feature $\hat{\mathbf{F}}_{t}$ and the reference feature $\mathbf{F}_{1}$, we first compute the 4D correlation volume between all pairs following optical flow literature [50]. The motion feature $f_{m}^{i}$ is obtained through a lightweight motion encoder, taking current estimation $\mathbf{f}_{1\rightarrow t}^{i}$, $\mathbf{v}_{1\rightarrow t}^{i}$, and correlation volume as input. Finally, we produce an update direction for flow and visibility with a GRU unit as in [33]:
@@ 3.4 Visibility-guided Splatting
To achieve this, forward warping via splatting [40] becomes a natural choice. Specifically, for every position $(x_{1},y_{1})$, we calculate its position in the target frame by adding the optical flow. Then bilinear kernel $b(\cdot)$ is employed for weighted summation as follows:
@@
$F_{t}^{v}=\frac{\sum\limits^{\rightarrow}(\mathbf{v}_{1\rightarrow t}\cdot\mathbf{F}_{1},\mathbf{f}_{1\rightarrow t})}{\sum\limits^{\rightarrow}(\mathbf{v}_{1\rightarrow t},\mathbf{f}_{1\rightarrow t})},$ (7)
@@
Lastly, the memory bank maintains two FIFO queues of keys and values, caching a window of $L$ recent first-frame-derived features. We append the splatting results $F_{t}^{v}$ into the queue of memory value. As for the memory key, we reuse the query feature $q$ of $\mathbf{F}_{t}$ for simplicity.
@@ 4.2 TAP-Vid and RoboTAP
We further provide qualitative comparisons with Online TAPIR [51], which is currently the best method for online tracking, on challenging videos from DAVIS in Fig. 5. SPOT outperforms Online TAPIR in handling complex occlusion, estimating better and consistent motion state of the car and people behind distractors (the first and second examples of Fig. 5), achieves better tracking during multi-subjects interaction (the third example), and effectively tracks small part of the animal (the fourth example).
@@ 4.4 Ablation study
Discussion of Memory Bank Length. We have demonstrated empirically that online dense point tracking can be resolved effectively by the memory mechanism. The intuition is that discriminative information can be propagated from the first frame to the latest frame within memory via already predicted accurate optical flow effectively, and this information can be readout by the attention mechanism based on feature similarity. As there is typically one or two timesteps between the current frame and the frame within memory, the feature similarity is accurate enough for information retrieval. This two-step information propagation mechanism is the cornerstone of the success of our approach. However, the success of two-step information propagation depends on the success of each step. Therefore, for accurate attention-based readout without suffering from the drifting problem, the length of the memory bank is typically small (we set the upper bound of the frame number to 3 in SPOT as verified in Tab. 4). If there is a video with a long time of occlusion, then our memory bank may not provide useful information due to occlusion and SPOT may fail to track the points. To deal with such scenarios, one possible future work is to introduce persistent object-level memory, for the objects within the first frame, that updates adaptively and survives through the long occlusion.

### v2

@@ Abstract
Dense point tracking is a challenging task requiring the continuous tracking of every point in the initial frame throughout a substantial portion of a video, even in the presence of occlusions. Traditional methods use optical flow models to directly estimate long-range motion, but they often suffer from appearance drifting without considering temporal consistency. Recent point tracking algorithms usually depend on sliding windows for indirect information propagation from the first frame to the current one, which is slow and less effective for long-range tracking. To account for temporal consistency and enable efficient information propagation, we present a lightweight and fast model with Streaming memory for dense POint Tracking and online video processing. The SPOT framework features three core components: a customized memory reading module for feature enhancement, a sensory memory for short-term motion dynamics modeling, and a visibility-guided splatting module for accurate information propagation. This combination enables SPOT to perform dense point tracking with state-of-the-art accuracy on the CVO benchmark, as well as comparable or superior performance to offline models on sparse tracking benchmarks such as TAP-Vid and RoboTAP. Notably, SPOT with 10$\times$ smaller parameter numbers operates at least 2$\times$ faster than previous state-of-the-art models while maintaining the best performance on CVO. We will release the models and codes at: https://dqiaole.github.io/SPOT/.
@@ 3.2 Feature Enhancement with Memory Reading
Specifically, the original feature $\mathbf{F}_{t}$ is projected through a linear projection layer for generating the query feature $q$. Then a dot-product similarity calculation is performed against the memory key followed by a softmax for normalization. This normalized score is used to aggregate the memory value to obtain the readout features $\mathbf{M}_{t}$. The process can be formulated as:
$\mathbf{M}_{t}=\mathrm{Softmax}(1/\sqrt{D_{k}}\times q\times k^{T})\times v,$ (1)
@@
Yet, we would like to point out that it is inappropriate to use the readout features $\mathbf{M}_{t}$ for flow regression directly. As our memory value is obtained through splatting the reference feature $\mathbf{F}_{1}$ with the previously predicted flow (which will be introduced in Sec. 3.4), there are many blank hole artifacts [1, 33] due to disocclusion [29]. These disocclusion artifacts will hamper the network training and degrade the performance of flow regression. Inspired by the typically used inpainting technique for hole filling [1], we further propose a fusion layer for artifacts remedy. Fortunately, we find that one single convolution layer performs well enough, and we “inpaint” the feature through the original feature $\mathbf{F}_{t}$:
${\mathbf{E}}_{t}=\mathbf{F}_{t}+\mathrm{Conv}(\mathbf{F}_{t}\oplus\mathbf{M}_{t}),$ (2)
where $\oplus$ denotes concatenation and ${\mathbf{E}}_{t}$ is the final enhanced feature that is ready for the next stage of flow regression.
@@ 3.3 Flow Decoding and Sensory Memory
Figure 3: Overview of optical flow decoder within Fig. 2. Motion feature is obtained through flow lookup. We update the flow with sensory memory, and update the sensory memory with the final motion feature. For the sake of simplicity, visibility prediction and context feature are omitted here.
With the memory-enhanced feature ${\mathbf{E}}_{t}$ and the reference feature $\mathbf{F}_{1}$, we first compute the 4D correlation volume between all pairs following optical flow literature [54]. The motion feature $f_{m}^{i}$ is obtained through a lightweight motion encoder, taking current estimation $\mathbf{f}_{1\rightarrow t}^{i}$, $\mathbf{v}_{1\rightarrow t}^{i}$, and correlation volume as input. Finally, we produce an update direction for flow and visibility with a GRU unit as in [35]:
@@ 3.4 Visibility-guided Splatting
To achieve this, forward warping via splatting [43] becomes a natural choice. Specifically, for every position $(x_{1},y_{1})$, we calculate its position in the target frame by adding the optical flow. Then bilinear kernel $b(\cdot)$ (details in Supp.) is employed for weighted summation as follows:
@@
$F_{t}^{1\rightarrow t}=\frac{\sum\limits^{\rightarrow}(\mathbf{v}_{1\rightarrow t}\cdot\mathbf{F}_{1},\mathbf{f}_{1\rightarrow t})}{\sum\limits^{\rightarrow}(\mathbf{v}_{1\rightarrow t},\mathbf{f}_{1\rightarrow t})},$ (7)
@@
Lastly, the memory bank maintains two FIFO queues of keys and values, caching a window of $L$ recent first-frame-derived features. We append the splatting results $F_{t}^{1\rightarrow t}$ into the queue of memory value. As for the memory key, we reuse the query feature $q$ of $\mathbf{F}_{t}$ for simplicity.
@@ 4.2 TAP-Vid and RoboTAP
We further provide qualitative comparisons with Online TAPIR [55], which is currently the best method for online tracking, on challenging videos from DAVIS in Fig. 5. SPOT outperforms Online TAPIR in handling complex occlusion, estimating better and consistent motion state of the car and people behind distractors (the first and second examples of Fig. 5), achieves better tracking during multi-subjects interaction (the third example), and effectively tracks small part of the animal (the fourth example). Animated video results can be found in our project page.
@@ 4.4 Ablation study
Discussion. We have demonstrated empirically that online dense point tracking can be resolved effectively by the memory mechanism. The intuition is that discriminative information can be propagated from the first frame to the latest frame within memory via already predicted accurate optical flow effectively, and this information can be readout by the attention mechanism based on feature similarity. As there is typically one or two timesteps between the current frame and the frame within memory, the feature similarity is accurate enough for information retrieval. This two-step information propagation mechanism is the cornerstone of the success of our approach.
Limitation. However, the success of two-step propagation depends on the success of each step. Therefore, for accurate attention-based readout without suffering from the drifting problem, the length of the memory bank is typically small (3 in SPOT as verified in Tab. 4). If there is a video with a extreme long time of occlusion, then SPOT may degrades to pairwise method, which is the main limitation of SPOT. To deal with such scenarios, one possible future work is to introduce persistent object-level memory, for the objects within the first frame, that updates adaptively and survives through the long occlusion. Besides, our SPOT is also not robust to extreme fast motion and texture-less region. And we provide more failure cases in Supplementary.
@@ 5 Conclusion
Acknowledgements
This work was sponsored by Doubao Fund, ByteDance. Yanwei Fu is also with Shanghai Innovation Institute, Institute of Trustworthy Embodied AI, Fudan University, and ISTBI–ZJNU Algorithm Centre for Brain-inspired Intelligence, Zhejiang Normal University.
Supplementary Material
6 Implementation Details
Network Details. We adopt the classical optical flow estimation network RAFT [54] as our backbone. Besides, following the modification of DOT [35] to the architecture of RAFT, we also use a stride 1 in the first convolutional layer of the image encoder and predict the visibility mask by an additional mask decoder. The channel number of different features is set as follows: $D$ of encoded feature from the image encoder is 256, $D_{k}$ of memory key is 128, $D_{v}$ of memory value is 256, and $D_{s}$ of sensory memory is 128. In addition, the bilinear kernel $b(\cdot)$ used in our visibility-guided splatting has the following formulation:
$b(\Delta)=\max(0,1-|\Delta_{x}|)\cdot\max(0,1-|\Delta_{y}|),$ (8)
where $\Delta=(x_{1},y_{1})+\mathbf{f}_{1\rightarrow t}^{N}\left[(x_{1},y_{1})\right]-(x_{t},y_{t})$.
Loss Function. The loss function of SPOT consists of the $l_{1}$ loss [54] for the predicted flows and binary cross-entropy loss for the visibility prediction. Specifically, we use exponentially increasing weights for predictions from different GRU iterations. Given ground-truth optical flow $\mathbf{f}_{1\rightarrow t}^{gt}$ and visibility mask $\mathbf{v}_{1\rightarrow t}^{gt}$, our loss function is defined as:
$\sum_{i=1}^{N}0.8^{N-i}\left[\lambda||\mathbf{f}_{1\rightarrow t}^{gt}-\mathbf{f}_{1\rightarrow t}^{i}||_{1}+\mathrm{BCE}(\mathbf{v}_{1\rightarrow t}^{gt},\mathbf{v}_{1\rightarrow t}^{i})\right],$
where $\lambda$ is set to 1000 empirically.
Training Details. We employ FlashAttention-2 [14] for fast attention computation within the memory reading module. During training, we use Adam optimizer with one-cycle [50] learning rate on eight NVIDIA H100 GPUs. The learning rate is 1e-4. The batch size is 24 for the first training stage (i.e., 500k steps on optical flow dataset Kubric-CVO [57]) and 8 for the second one (i.e., 100k steps on point tracking dataset Kubric-MOVi-F [23]).
Evaluation Details. We set the iteration number $N$ of flow decoding to 16 by default, as 16 iterations already achieve the peak accuracy on real-world videos from DAVIS [46]. In contrast, we set it to 32 on Kinetics [7], RGB-Stacking [36], and RoboTAP [55]. Because we find that more iterations of GRU can improve the accuracy of these three datasets. These may be due to the different motion characteristics (e.g., faster motion of human actions and weak texture of robotic scenes) of these three datasets.
7 More Quantitative Analysis
Online Setup under Original Window Size. We additionally give a setup without modifying the window size of existing offline models: when processing frame $t$, we provide the model with all prior frames. This process is repeated for each frame sequentially, ensuring predictions are based solely on past frames. This gives an upper bound of performance for reference though with extremely slow speed. We use $\ddagger$ to denote this setup. Due to the extremely slow inference speed of this setup, we only provide the result on DAVIS (First) in Tab. 5: SPOT still beats other online versions by a large margin, with a much faster inference speed. It still supports our contributions of strong performance and superior efficiency.
Table 5: Point tracking results on DAVIS (First). $\ddagger$ represents evaluated in an online fashion under the original window size with extremely slow speed.
Online AJ $\uparrow$ $<\delta_{avg}^{x}\uparrow$ OA $\uparrow$
TAPIR 56.7 70.2 85.7
CoTracker2 55.9 68.7 83.7
SpatialTracker 57.3 70.6 85.0
DOT 57.3 69.7 85.2
Online TAPIR 56.2 69.3 84.6
Ours 61.5 75.0 88.9
Further Discussion on Forward Splatting and Backward Warping. The main idea of SPOT is breaking down long-term tracking into two simpler steps, i.e., long-range propagation with flow and similarity-based short-range retrieval. SPOT instantiates propagation with splatting and retrieval with attention. We can also first use attention to algin $\mathbf{F}_{t}$ to memory frames, then backward warp them to frame 1, i.e., a reversed pipeline of SPOT. Though there are many-to-one mapping and empty regions in splatting, we tackle them with visibility mask and inpainting. However, there are also problems with backward warping, i.e., warping error foreground for occluded region, warping out-of-frame empty regions. These problems are caused by underlying physical motion. Therefore, there is no preference for which long-range propagation method to choose. The important point is the framework of two-step breakdown. Here, we further provide an additional ablation experiment: we warp the features of memory frames to the first frame directly, without the stage of attention for short-range retrieval. As shown in Tab. 6, though the ablated model achieve similar results as our SPOT on short videos, it fails to generalize to longer videos (CVO Extended) due to error accumulation.
Table 6: Alation results on CVO dataset. We ablate the attention block of SPOT here.
Method CVO (Final) CVO (Extended)
EPE $\downarrow$ (all / vis / occ) EPE $\downarrow$ (all / vis / occ)
SPOT 1.17 / 0.67 / 3.49 6.42 / 3.86 / 9.98
Attention Ablation 1.18 / 0.69 / 3.50 395.35 / 379.44 / 428.48
Computational Complexity. SPOT only maintains 3 frames memory, the computational complexity will not increase after time step $t$ being larger than 3. In addition, we provide curve of GPU memory and speed w.r.t video resolution in Fig. 7. SPOT runs much faster than DOT across varying resolutions up to 1024x1024, while consumes similar memory. The main computational overhead still lies in RAFT. So, we believe recent progress in efficient and high-resolution optical flow estimation can greatly and immediately benefits SPOT due to the unified architecture design.
Figure 7: GPU memory and inference FPS v.s. input resolution.
Compared to Recent Methods. Track-On [2] and DELTA [42] are very recent works that focus on online sparse tracking and offline dense tracking, respectively. However, SPOT tackles online dense tracking. Here, we provide an efficiency comparison with them under our setting. On H100, SPOT tracks 512x512 videos (around 262k points) at 12.4 FPS with 4.15GB GPU memory. But Track-On can only tracks up to 21k points at 1.23 FPS with 77.2GB GPU memory; DELTA runs at 0.19 FPS with 49.58GB GPU memory. Considering the significant latency and memory consumption, both are not suitable for online dense point tracking we consider here.
Accuracy vs. Temporal Interval. We repurpose TAP-Vid to evaluate performance across varying temporal intervals explicitly. Specifically, we employ the DAVIS (First) split here and evaluate the performance for each temporal interval (up to 30 here) without averaging the results across the temporal intervals. Fig. 8 shows that AJ will degrade gradually as the temporal interval increases, while SPOT generally outperforms Online TAPIR [55] on all temporal intervals. Besides, AJ of SPOT w/o memory degrades rapidly, illustrating the important role of memory module. Though length of our memory bank is only 3, SPOT propagates information from first frame to memory frames directly, greatly alleviating the information degradation and helping handling long videos.
Figure 8: AJ of our SPOT, SPOT w/o Memory, and Online TAPIR across varying temporal intervals on DAVIS (First).
8 More Qualitative Results
Qualitative Analysis of Memory Bank. We provide qualitative ablation of memory bank on video with large appearance variations in Fig. 9. As shown in Fig. 9, removing the memory bank leads to the failure of tracking due to appearance changes. Our SPOT with memory bank can successfully track the points of cars and overcome the challenge of appearance variations.
Figure 9: Qualitative ablation of memory bank on challenging real-world video.
Qualitative Result on Long Occlusion. SPOT introduces visibility mask during splatting (Eq. 7 of main paper). And splatted feature of occluded region will be all zeros and no effective information can be read out through attention. So SPOT ‘degrades’ to pairwise method in such extreme case. We provide a such case in Fig. 10, where right woman (red color) is totally occluded by man for 10 frames (longer than 3 frames memory). Once the woman reappears, SPOT successfully locates the woman and recovers from occlusions.
Figure 10: Qualitative result on long occlusion.
Qualitative Results on CVO. We provide more qualitative comparison results with the previous state-of-the-art method for dense tracking, DOT [35], on long-range optical flow benchmark CVO in Fig. 11. The areas where our SPOT achieves substantial improvements are highlighted with bounding boxes. Please zoom in for more details. DOT typically fails to estimate the motion of small objects, occluded objects, and objects with weak textures. By contrast, our SPOT successfully estimates the motion for these hard cases.
Qualitative Results on Real-world Videos. We also provide more qualitative results with the previous state-of-the-art method for online tracking, Online TAPIR [55], on real-world videos from DAVIS in Fig. 12. Fig. 12 shows that our SPOT has superior performance on real-world videos. Please zoom in for more details.
Figure 11: More qualitative comparison on the CVO (Extended). Notable areas are marked by a bounding box. Please zoom in for details.
Figure 12: More qualitative comparison on DAVIS. For each sequence, we show tracking results of Online TAPIR [55] and SPOT. Only foreground points of the first frame are visualized, each point is displayed with a different color and overlayed with white stripes if occluded. Please zoom in for details.
Failure Cases. We provide some failure cases in Fig. 13. Our SPOT fails to track the bike after extreme long occlusion, i.e., more than 20 occluded frames in the first case. Besides, SPOT cannot distinguish texture-less objects properly, especially there are four similar fast-moving ducks in the second case. Finally, for fast motion shown in the third and fourth cases, SPOT may lose the track of thin object or even the whole object.
Figure 13: Failure cases on DAVIS. Please zoom in for details.